# Read Between the Layers: Leveraging Intra-Layer Representations for Rehearsal-Free Continual Learning with Pre-Trained Models

### v1

@@ 4.2 Combination with parameter-efficient model adaptation
A benefit of class-prototype methods for CL is that they can be orthogonally combined with adaptation techniques to refine the pre-trained representations. As accumulated class prototypes are subject to discrepancy upon distributional shifts during supervised fine-tuning, any adaptation strategies to the latent representations should be carefully tailored to the class-prototype method used. Previous works [44, 72] have found adaptation to downstream continual tasks during First Session Adaptation (FSA) on $\mathcal{D}_{1}$ as sufficient to bridge the domain gap while maintaining full compatibility with CL.
To maintain the powerful generic features of the pre-trained backbone, we choose to apply FSA to additional parameter-efficient transfer learning (PETL) parameters while keeping the backbone frozen throughout. To update PETL parameters during FSA, we train a linear classification head with $N_{1}$ outputs via Adam optimizer and cross-entropy loss and discard the classification head afterward. During the CL phase, we keep all model parameters frozen and update ${\bm{G}}$ and ${\bm{c}}_{y}$ only.
We follow prior works [72, 40] and experiment with Visual Prompt Tuning (VPT) [22], Scaling and Shifting of Features (SSF) [34], and AdaptFormer [5] as PETL methods and refer to the cited papers for details. The pseudocode of the LayUP algorithm for the CIL setting is depicted in Alg. 1. For the OCL setting, the FSA stage is omitted and a fixed $\lambda$ is used, as a dynamic $\lambda$ search requires multiple passes over the input data.
@@
4.3 Memory and runtime complexity comparisons
We compare the memory and runtime requirements of LayUP with the three competitive CL methods ADAM [72], SLCA [69], and RanPAC [40] for a typical class count of $C=200$. We neglect the memory cost of the ViT backbone, the fully connected classification layer, and the class prototypes, as they are similar to all baselines. LayUP stores $\sim$1M PETL parameters (AdaptFormer [5]) and an additional intra-layer Gram matrix in memory. The size of this layer depends on the choice of $k$ and has $(d_{L-k+1}+\dots+d_{L-1}+d_{L})^{2}$ entries (note that for ViT-B/16 [12] model, every layer outputs tokens of the same dimensionality, such that $d_{l}=768$ for $l\in\{1,\dots,L\}$). For $k=6$, which we use in our final experiments, the matrix stores $\sim$21M values. RanPAC updates a Gram matrix of size 100M and requires an additional $\sim$11M parameters for the random projection layer and PETL parameters. SLCA stores $C$ additional covariance matrices of size $d_{L}^{2}$ with a total of $\sim$118M entries. Finally, ADAM stores a second copy of the feature extractor, necessitating 84M additional parameters. Assuming a comparable memory cost for model parameters and matrix entries, our method reduces additional memory requirements by $81\%$, $82\%$, and $75\%$ compared with RanPAC, SLCA, and ADAM, respectively.
[…]
5 Experiments
In what follows, we conduct a series of experiments to investigate the performance of our approach under different CL settings. First, we provide an overview of the datasets and implementation details in Sec. 5.1. Second, we perform empirical comparisons of our method with strong recent baselines under the CIL and the more challenging OCL settings in Sec. 5.2 and Sec. 5.3, respectively. Third, we conduct ablation studies in Sec. 5.4. We conclude our empirical analysis by investigating the contribution of different choices of $k$ last layers to the classifier performance in Sec. 5.5.
5.1 Datasets and implementation details
Method CIFAR IN-R IN-A CUB OB VTAB Cars
Joint full FT $93.6$ $86.6$ $71.0$ $91.1$ $80.3$ $95.5$ $83.7$
Joint linear probe $87.9$ $71.2$ $56.4$ $89.1$ $78.8$ $90.4$ $66.4$
L2P [65] $84.6$ $72.5$ $42.5$ $65.2$ $64.7$ $77.1$ $38.2$
DualPrompt [64] $81.3$ $71.0$ $45.4$ $68.5$ $65.5$ $81.2$ $40.1$
CODA-P [56] $86.3$ $75.5$
ADAM [72] $87.6$ $72.3$ $52.6$ $87.1$ $74.3$ $84.3$ $41.4$
SLCA [69] $91.5$ $77.0$ $59.8^{*}$ $84.7$ $73.1^{*}$ $89.2^{*}$ $67.7$
RanPAC [40] $78.1$ $61.8$ $92.6$ $77.7$
LayUP $91.5$ $88.1$ $77.6$
Ablations
w/o FSA $88.7$ $73.4$ $51.6$ $86.9$ $77.0$ $92.9$ $77.5$
${k=1}$ $90.4$ $79.0$ $61.5$ $86.0$ $74.3$ $93.1$ $75.5$
$k=1$, w/o FSA $86.5$ $69.2$ $55.6$ $85.4$ $72.8$ $92.2$ $69.2$
LayNMC $88.1$ $75.2$ $56.2$ $85.5$ $71.8$ $87.5$ $57.6$
Table 1: Average accuracy (%) after training: Comparison of prompting, backbone fine-tuning, and class-prototype methods for the CIL setting. Results are taken from [40] except results for SLCA and Coda-P, which are taken from the respective papers. Results for SLCA marked with ($*{}^{*}$) were reproduced using the officially released code. The remaining results for CODA-P could not be reproduced due to excessive GPU memory requirements for training ($>$24GB).
Following prior works [64, 65, 72, 69, 40], we experiment with two ViT-B/16 [12] models, the former with self-supervised pre-training on ImageNet-21K (ViT-B/16-IN21K) [52], the latter with additional supervised fine-tuning on ImageNet-1K (ViT-B/16-IN1K) [30]. During adaptation in the CIL setting, PETL parameters are trained using a batch size of 48 for 20 epochs and Adam [26] with momentum for optimization. We employ learning rate scheduling via cosine annealing, starting with 0.03. We train five prompt tokens for VPT and use a bottleneck dimension of 16 for AdaptFormer. All baselines use the same aforementioned ViT-B/16 backbones and are trained on seven representative split datasets CIFAR-100 (CIFAR) [29], ImageNet-R (IN-R) [16], ImageNet-A (IN-A) [17], CUB-200 (CUB) [60], OmniBenchmark (OB) [70], Visual Task Adaptation Benchmark (VTAB) [68], and Stanford Cars-196 (Cars) [28]. None of the methods compared in Tab. 1 and Tab. 2 require a data buffer for rehearsal. We use $T=10$ for all datasets, except $T=5$ for VTAB, which is a benchmark composed of five different datasets to be learned consecutively. (We report additional results for $T=5$ and $T=20$ in Sec. C.4.) VTAB can be considered as a hybrid class- and domain-incremental learning benchmark according to the definition in [58], as label spaces of each task are partially overlapping. Dataset statistics can be found in Appendix B. For comparison, we use the average accuracy [36] metric $A_{t}=\frac{1}{t}\sum_{i=1}^{t}R_{t,i}$, with $R_{t,i}$ denoting the classification accuracy on the $i$th task after training on the $t$th task. All experiments are conducted with $k=6$ unless stated otherwise. Final accuracy $A_{T}$ after learning the last task is reported for random seed 1993 and the best combination of PETL method and ViT backbone (similar to [40, 72]) in the main paper. We refer to Sec. C.1 for analysis of each $A_{t}$, forgetting measures, and variability across random initialization, and to Sec. C.3 for average accuracy over training for different PETL methods and pre-trained backbones in the CIL setting.
5.2 Performance in the CIL setting
Method CIFAR IN-R IN-A CUB OB VTAB Cars
Sequential FT $12.8$ $5.3$ $6.9$ $5.1$ $3.3$ $7.3$ $10.9$
NMC [72] $83.4$ $61.2$ $49.3$ $86.7$ $73.2$ $88.4$ $37.9$
@@ 5.2 Performance in the CIL setting
LayNMC $80.8$ $59.7$ $50.4$ $82.9$ $69.3$ $85.9$ $40.0$
Table 2: Average accuracy (%) after training: Comparison of class-prototype methods for the OCL setting. Results for RanPAC and NMC are reproduced according to the officially released code in [40].
We first compare LayUP with several prompt learning, fine-tuning, and class-prototype methods for CL. We additionally report results for joint training of a randomly initialized linear probe and joint full fine-tuning (FT) of the ViT-B/16 backbone. Final accuracy scores are shown in Tab. 1.
LayUP surpasses all baselines for four of the seven split datasets. Interestingly, these four datasets (IN-R, IN-A, VTAB, Cars) have two distinctive characteristics: First, they have the highest domain gap (indicated by the cosine distance to the pre-trained miniImageNet domain for the offline setting and 50 shots, cf. [44]) among all datasets. Second, they have significantly less training data than CIFAR, OB, and CUB (cf. Appendix B). The former confirms our initial hypothesis that intra-layer representations are more domain-invariant and consequently more robust to large distributional shifts. The latter indicates that especially in the low-data regime, in contrast to other methods that tend to overfit to the target domain, LayUP constructs class prototypes and decision boundaries that generalize well even if the amount of training data is scarce compared with the data used for ViT pre-training. It is further noteworthy that RanPAC, which is the only baseline that LayUP does not consistently outperform, is considerably more expensive regarding both memory and computation (cf. Sec. 4.3).
5.3 Performance in the OCL setting
We are interested in assessing the performance of our method in the challenging OCL setting, where only a single pass over the continual data stream is allowed. All baselines are class-prototype methods for continual learning on a frozen embedding network (thus we omit FSA stages for RanPAC, and our approach), except for sequential fine-tuning, where we update the parameters of the pre-trained ViT via cross-entropy loss and Adam [26] optimizer during a single epoch. As the choice of ridge regression [18] parameter as used in RanPAC and our method requires prior knowledge about the downstream continual data–which we do not have in a realistic streaming learning setting–we compare with two simplified versions of the Gram matrix inversion $({\bm{G}}+\lambda{\bm{I}})^{-1}$ (cf. Eq. 5 and Eq. 2): In the first variant, we omit $\lambda$-based regularization completely, such that $\lambda=0$ and the Gram matrix is inverted without regularization (i.e., ${\bm{G}}^{-1}$). In the second variant, e.g., as used in [44], we compare with $\lambda=1$, such that we linearly transform identity-regularized class prototypes (i.e., $({\bm{G}}+{\bm{I}})^{-1}$).
As indicated in Tab. 2, unrestricted fine-tuning of the backbone is detrimental to the generalizability of the pre-trained embeddings and leads to forgetting and low performance. Consequently, the sequential FT baseline is outperformed by class-prototype methods in the OCL setting for most benchmarks. LayUP is largely robust to missing regularization ($\lambda=0$), with the exception of a drop in performance on VTAB. Moreover, it maintains a high performance across all benchmarks for regularization with $\lambda=1$. On the contrary, RanPAC shows high variability in performance across benchmarks under the absence of regularization and for regularization with $\lambda=1$, partly resulting in near-zero accuracy. Such results can be explained by the fact that the high-dimensional Gram matrix obtained after random projections in RanPAC overfits the training data and thus struggles to generalize, especially under a complete lack of regularization. The higher overall robustness to different OCL settings makes LayUP a favorable approach in continual learning scenarios where the length of the input stream and the nature of the data might not be known in advance.
@@ 5.4 Ablation Study
[…]
LayUP consistently benefits from intra-layer representations for class-prototype construction, with the sole exception of training on IN-A in the OCL setting (Tab. 2). It further benefits from leveraging second-order statistics for the decorrelation of class prototypes and first session training to bridge the domain gap to downstream CL tasks. Finally, it is the combination of (i), (ii), and (iii) that yields the highest performance and thus makes LayUP a strong and competitive class-prototype method for CL with pre-trained models.

### v2

@@ 4.2 Combination with parameter-efficient model adaptation
A benefit of class-prototype methods for CL is that they can be orthogonally combined with adaptation techniques to refine the pre-trained representations. As accumulated class prototypes are subject to discrepancy upon distributional shifts during supervised fine-tuning, any adaptation strategies to the latent representations should be carefully tailored to the class-prototype method used. Previous works (Panos et al., 2023; Zhou et al., 2023b) find adaptation to downstream continual tasks during First Session Adaptation (FSA) on $\mathcal{D}_{1}$ as sufficient to bridge the domain gap while maintaining full compatibility with CL.
@@
To maintain the powerful generic features of the pre-trained backbone, we choose to apply FSA to an additional set of learnable PETL parameters while keeping the backbone frozen throughout. To update PETL parameters during FSA, we train a linear classification head with $N_{1}$ outputs via Adam optimizer (Kingma & Ba, 2015) and cross-entropy loss only during the first CL stage and discard the classification head afterward. During the CL phase, we keep all model parameters frozen and update ${\bm{G}}$ and ${\bm{c}}_{y}$ only. We follow prior works (Zhou et al., 2023b; McDonnell et al., 2023) and experiment with Visual Prompt Tuning (VPT) (Jia et al., 2022), Scaling and Shifting of Features (SSF) (Lian et al., 2022), and AdaptFormer (Chen et al., 2022) as PETL methods and refer to the cited papers for details.
The pseudocode of the LayUP algorithm for the CIL setting is presented in Alg. 1. It is noteworthy that both ${\bm{G}}$ and ${\bm{c}}_{y_{t}}$ can be updated incrementally, one sample at a time, and Eq. 5 can be applied after every sample, thereby ensuring full compatibility with online (or, streaming) learning settings. For the OCL comparison, the FSA stage is omitted, and a fixed $\lambda$ is utilized, since a dynamic search for $\lambda$ necessitates multiple passes over the input data.
4.3 Memory and runtime complexity comparisons
@@
[…]
5 Experiments
In what follows, we conduct a series of experiments to assess our approach under various CL settings. We start with an overview of the datasets and implementation details in Sec. 5.1, followed by empirical comparisons of our method against recent strong baselines under CIL, DIL, and OCL settings in Sec. 5.2 and Sec. 5.3. Ablation studies are presented in Sec. 5.4. We explore the impact of different configurations of the last $k$ layers on classifier performance in Sec. 5.5 and conclude our empirical analysis by investigating the range of classes that benefit from intra-layer features in Sec. 5.6.
5.1 Datasets and implementation details
Following prior works (Wang et al., 2022b; c; Zhou et al., 2023b; Zhang et al., 2023; McDonnell et al., 2023), we experiment with two ViT-B/16 (Dosovitskiy et al., 2021) models, the former with self-supervised pre-training on ImageNet-21K (ViT-B/16-IN21K) (Ridnik et al., 2021), the latter with additional supervised fine-tuning on ImageNet-1K (ViT-B/16-IN1K) (Krizhevsky et al., 2012). During adaptation in the CIL and DIL settings, PETL parameters are trained using a batch size of 48 for 20 epochs, Adam (Kingma & Ba, 2015) with momentum for optimization, and learning rate scheduling via cosine annealing, starting with 0.03. We train five prompt tokens for VPT and use a bottleneck dimension of 16 for AdaptFormer. All baselines are rehearsal-free and trained on the same aforementioned ViT backbones.
For the CIL and OCL settings, the seven representative split datasets are CIFAR-100 (CIFAR) (Krizhevsky, 2009), ImageNet-R (IN-R) (Hendrycks et al., 2021a), ImageNet-A (IN-A) (Hendrycks et al., 2021b), CUB-200 (CUB) (Wah et al., 2011), OmniBenchmark (OB) (Zhang et al., 2022), Visual Task Adaptation Benchmark (VTAB) (Zhai et al., 2020), and Stanford Cars-196 (Cars) (Krause et al., 2013). We use $T=10$ for all datasets, except VTAB, which has a commonly used task count of $T=5$. Additional results for task counts $T=5$ and $T=20$ can be found in Sec. C.4. For the DIL setting, baselines are trained on Continual Deepfake Detection Benchmark Hard (CDDB-H) (Li et al., 2023) with $T=5$, a sub-sampled version of DomainNet (S-DomainNet) (Peng et al., 2019) with $T=6$, and the domain-incremental formulation of ImageNet-R (IN-R (D)) with $T=15$. Detailed descriptions and summary statistics of all datasets can be found in Appendix B.
For comparison, we use the average accuracy (Lopez-Paz & Ranzato, 2017) metric $A_{t}=\frac{1}{t}\sum_{i=1}^{t}R_{t,i}$, with $R_{t,i}$ denoting the classification accuracy on the $i$th task after training on the $t$th task. Based on the observations detailed in Sec. 4.1, utilizing features from the second half of the network layers for prototype construction emerges as an effective yet cost-efficient strategy in terms of memory and computational resources. Consequently, all experiments are conducted with $k=6$, unless specified otherwise. For a comprehensive comparison of performance across different choices of $k$ for all datasets, refer to Sec. 5.5. In all experiments in the main paper, the reported metric is average accuracy $A_{T}$ after learning the last task for random seed 1993 and the best combination of PETL method and ViT backbone (similar to McDonnell et al. (2023); Zhou et al. (2023b)). We refer to Sec. C.1 for analysis of each $A_{t}$, forgetting measures, and variability across random initialization, and to Sec. C.3 for average accuracy over training for different PETL methods and pre-trained backbones in the CIL setting.
5.2 Performance in the CIL and DIL settings
Method CIFAR IN-R IN-A CUB OB VTAB Cars
Joint FT (full) 93.6 86.6 71.0 91.1 80.3 92.5 83.7
Joint FT (linear probe) 87.9 71.2 56.4 89.1 78.8 90.4 66.4
L2P (Wang et al., 2022c) 84.6 72.5 42.5 65.2 64.7 77.1 38.2
DualPrompt (Wang et al., 2022b) 81.3 71.0 45.4 68.5 65.5 81.2 40.1
CODA-P (Smith et al., 2023a) 86.3 75.5 74.5 79.5 68.7 87.4 43.2
NMC+FSA (Janson et al., 2022) 87.8 70.1 49.7 85.4 73.4 88.2 40.5
ADAM (Zhou et al., 2023b) 87.6 72.3 52.6 87.1 74.3 84.3 41.4
SLCA (Zhang et al., 2023) 91.5 77.0 59.8 84.7 73.1 89.2 67.7
RanPAC (McDonnell et al., 2023) 78.1 61.8 92.6 77.7
LayUP 92.0 88.1 77.6
Ablations
${k=1}$ 90.8 78.8 60.4 86.7 72.0 92.4 74.9
w/o FSA 88.7 73.1 57.7 86.2 77.0 92.6 78.6
$k=1$, w/o FSA 86.5 69.4 55.4 85.4 70.7 91.6 69.1
LayNMC 88.1 59.3 50.0 82.5 68.9 86.5 40.0
NMC 83.4 61.2 49.3 85.1 73.1 88.4 37.7
Table 1: Comparison of prompting, backbone fine-tuning, and class-prototype methods for the CIL setting. Results are taken from McDonnell et al. (2023) except results for SLCA marked with (), which are reproduced using the officially released code.
Method CDDB-H S-DomainNet IN-R (D)
Joint FT (full) 92.9 62.8 86.6
Joint FT (linear probe) 74.0 57.2 71.2
NMC+FSA (Janson et al., 2022) 49.8 44.0 76.3
ADAM (Zhou et al., 2023b) 70.7
RanPAC (McDonnell et al., 2023) 86.2 58.8 77.2
LayUP
Ablations
${k=1}$ 84.0 54.2 77.8
w/o FSA 86.7 58.3 73.3
$k=1$, w/o FSA 77.7 53.7 69.0
LayNMC 61.4 34.4 59.1
NMC 57.6 43.5 64.1
Table 2: Comparison of class-prototype methods for the DIL setting. Results for ADAM and RanPAC on CDDB-H are taken from McDonnell et al. (2023).
In the CIL setting, we compare LayUP with several prompt learning, fine-tuning, and class-prototype methods for CL. We additionally report results for joint fine-tuning (FT) of the full backbone and a linear probe. As shown in Tab. 1, LayUP surpasses all baselines for four of the seven split datasets. Interestingly, these four datasets (IN-R, IN-A, VTAB, Cars) have two distinctive characteristics: First, they have the highest domain gap among all datasets, as indicated by the cosine distance to the pre-trained miniImageNet domain for the offline setting and 50 shots (cf. Panos et al. (2023)). Second, they have significantly less training data than CIFAR, OB, and CUB (cf. Appendix B). The former confirms our initial hypothesis that intra-layer representations are more domain-invariant and consequently more robust to large distributional shifts from the source to the target domain. The latter indicates that especially in the low-data regime, in contrast to other methods that tend to overfit to the target domain, LayUP constructs class prototypes and decision boundaries that generalize well even if the amount of training data is scarce compared with the data used for ViT pre-training. It is further noteworthy that RanPAC, which is the only baseline that LayUP does not consistently outperform, is considerably more expensive regarding both memory and computation (cf. Sec. 4.3).
Results for the DIL setting are reported in Tab. 2. We compare our method with several strong class-prototype methods for CL and, similar to the CIL comparison, we additionally report results for joint fine-tuning of the backbone and a linear probe. LayUP consistently outperforms all baselines. These results confirm that our method effectively utilizes the domain-invariant information contained in the low- and mid-level intermediate features of the pre-trained model, thereby enhancing its robustness to domain shifts.
5.3 Performance in the OCL setting
We are interested in assessing the performance of our method in the challenging OCL setting, where only a single pass over the continual data stream is allowed. All baselines are class-prototype methods for continual learning on a frozen embedding network (thus we omit FSA stages for NMC, RanPAC, and our approach), except for sequential fine-tuning, where we update the parameters of the pre-trained ViT via cross-entropy loss and Adam optimizer during a single epoch. Note that NMC exhibits identical behavior in both CIL and OCL settings, as it uses each training sample only once for running-mean calculation, consequently producing the same outcomes as the ablation baseline in Tab. 1. Given that the choice of the ridge regression parameter $\lambda$, as utilized in RanPAC and our method, necessitates prior knowledge about the downstream continual data—a requirement unmet in realistic streaming learning settings—we opt for comparison with two simplified versions of Gram matrix inversion $({\bm{G}}+\lambda{\bm{I}})^{-1}$ (cf. Eq. 2 and Eq. 5): In the first variant, we omit $\lambda$-based regularization completely, such that $\lambda=0$ and the Gram matrix is inverted without regularization (i.e., ${\bm{G}}^{-1}$). In the second variant, e.g., as used in Panos et al. (2023), we choose $\lambda=1$, such that we linearly transform identity-regularized class prototypes (i.e., $({\bm{G}}+{\bm{I}})^{-1}$).
Method CIFAR IN-R IN-A CUB OB VTAB Cars
Sequential FT 12.8 5.3 6.9 5.1 3.3 7.3 10.9
NMC (Janson et al., 2022) 83.4 61.2 49.3 85.1 73.1 88.4 37.7
RanPACλ=0 (McDonnell et al., 2023) 70.6 1.4 0.9 76.9 9.3 0.6
RanPACλ=1 (McDonnell et al., 2023) 70.6 0.9 0.7 76.9 9.0 0.6
LayUPλ=0 40.9 86.4 10.4 66.3
LayUPλ=1 51.6
@@ 5.3 Performance in the OCL setting
Table 3: Comparison of class-prototype methods for the OCL setting. Results for RanPAC and NMC are reproduced according to the officially released code in McDonnell et al. (2023).
As indicated in Tab. 3, unrestricted fine-tuning of the backbone is detrimental to the generalizability of the pre-trained embeddings and leads to forgetting and low performance. Consequently, the sequential FT baseline is outperformed by class-prototype methods in the OCL setting for most benchmarks. With the exception of VTAB, LayUP is largely robust to missing regularization ($\lambda=0$) and maintains a high performance across all benchmarks for regularization with $\lambda=1$. On the contrary, RanPAC exhibits high variability in performance across benchmarks, occasionally resulting in near-zero accuracy. This outcome is likely due to the high-dimensional prototypes and Gram matrix in RanPAC, obtained after random projections, overfitting the training data, which consequently hampers their ability to generalize. LayUP’s superior robustness in the online learning setting renders it a more favorable approach for continual learning scenarios where the length of the input stream and the nature of the data might not be known in advance.
$k$ CIFAR IN-R IN-A CUB OB VTAB Cars
1 89.2 79.2 60.8 85.3 71.6 92.7 75.5
2 90.1 79.9 62.2 86.9 73.9 90.7 78.7
3 90.2 80.7 62.5 87.3 74.5 93.2 81.4
4 90.4 81.1 62.0 87.6 76.7 93.1 81.7
5 90.7 81.5 62.8 87.3 77.0 93.0 82.4
6 91.0 81.2 62.2 87.3 77.5 92.2 82.5
7 90.9 81.6 63.4 87.5 77.7 93.4 82.6
8 90.8 81.4 62.8 87.7 77.2 92.1 82.3
9 91.0 81.5 62.1 87.4 78.3 92.0 82.6
10 90.9 81.4 63.4 87.9 77.4 94.0 81.5
11 91.1 81.5 62.0 87.6 78.1 93.7 81.4
12 90.8 80.8 62.8 88.0 77.8 93.9 82.3
Table 4: Comparison of maximum representation depths. LayUP performance for different values of $k$ for a pre-trained Vit-B/16-IN1K and FSA with AdaptFormer. The 1st, 2nd, and 3rd highest scores and the 1st, 2nd, and 3rd lowest scores are highlighted.
@@ 5.4 Ablation study
The purpose of the following ablation study is to demonstrate the advantages of leveraging intra-layer representations over solely utilizing features from the final representation layer (i.e., $k=1$) for class-prototype construction in CL. As LayUP integrates intra-layer representations with second-order statistics via Gram matrix inversion as well as first-session adaptation in the CIL and DIL settings, we consider three configurations for ablating intra-layer representations: (i) in the standard setting with both FSA and second-order statistics (LayUP vs. $k=1$), (ii) excluding the FSA stage, i.e., omitting phase A in Alg. 1 (w/o FSA vs. $k=1$, w/o FSA), and (iii) excluding FSA and second-order statistics, which corresponds to traditional nearest-mean classifiers (LayNMC vs. NMC). Results for the baselines constructed for settings (i)-(iii) can be found in the bottom sections of Tab. 1 and Tab. 2. For (i), enriching the final layer with intermediate features consistently results in absolute performance gains ranging from 1.2% to 5.7% (CIL) and 2.2% to 5.2% (DIL). For (ii), except for the VTAB dataset in the CIL setting, leveraging intra-layer features yields absolute accuracy improvements of up to 9.5% (CIL) and 9.0% (DIL). Lastly, for (iii), LayNMC does not show superior performance to the NMC baseline. Such results indicate that the benefits of intermediate features diminish when not decorrelated via Gram matrix inversion, as they contain more lower-level noisy information than final layer representations.
In the lower section of Tab. 3, further comparisons are made between ablated versions of LayUP and prototypes using final representations ($k=1$) in the OCL setting for two selected values of the regularization parameter $\lambda$. LayUP generally outperforms its ablated versions, although it demonstrates decreased performance in low-data benchmarks such as IN-A and VTAB, particularly when omitting regularization ($\lambda=0$). This decline in performance may stem from the increased dimensionality of class prototypes and the Gram matrix due to the addition of intra-layer representations, thereby elevating the risk of overfitting. Consequently, the absence of regularization, combined with limited training data, may render intra-layer features detrimental. However, in any other case, they enhance classification accuracy by 0.6% to 8.3%.